The temperature at a point is given by T = xyz. Find the average temperature in the cube

with opposite corners at (0,0,0) and (2,2,2).


1
Expert's answer
2022-03-23T04:58:26-0400

"T=xyz\\
(0,0,0), (2,2,2)\\
\frac{1}{2\cdot2\cdot2}
\int_0^2\int_0^2\int_0^2xyzdxdydz=\frac{1}{8}\int_0^2\int_0^2xy\frac{z^2}{2}|_0^2dxdy=\\
=\frac{1}{8}\int_0^2\int_0^22xydxdy=\frac{1}{4}\int_0^2x\frac{y^2}{2}|_0^2dx=\frac{1}{4}\int_0^22xdx=\\
=\frac{1}{2}\frac{x^2}{2}|_0^2=\frac{1}{2}\cdot2=1"